### v1

# AllSpark: a multimodal spatiotemporal general model

Abstract
RGB, infrared, multispectral, and other modal data fundamentally represent different observational approaches to the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, for a long time, due to the high heterogeneity in structure and semantics among various spatiotemporal modal data, the joint interpretation of multimodal spatiotemporal data has been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities, and this trade-off exhibits a progressively nonlinear nature as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model, aiming to strike a trade-off between the cohesion and autonomy among different modalities. Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, called AllSpark. Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities. To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data. To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatiotemporal modalities. And modal bridge is employed to achieve dimensional projection from each modality to the language modality. Finally, observing a gap between the model’s interpretation and downstream tasks, we designed task heads to enhance the model’s generalization capability on specific downstream tasks. Experiments indicate that AllSpark, without expert knowledge of the most spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model. This approach contributes to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
keywords:
Multimodal, Language, Spatiotemporal, General Model
1 Introduction
Benefiting from the increasingly diverse observational methods for spatiotemporal scenes, geographic objects can be described using various spatiotemporal modalities such as RGB, infrared, SAR, multispectral, graphs, and spatiotemporal trajectories. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
However, due to the inherent differences in the mechanisms of each modality, various modalities are often highly heterogeneous in both structure and semantics. For example, in terms of structure, a table is composed of rows and columns. Point clouds are represented by three-dimensional coordinates along with various feature values. Text is constituted by sequences of words. In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
For a long time, constrained by the highly heterogeneity across various modalities as mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNetqi2017pointnet for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For the language modality, Vaswani et al. introduced Transformervaswani2017attention , focusing on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed GCNkipf2016semi , based on the adjacency relationships between nodes in a graph. In multimodal research, fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensingkulkarni2020pixel . Text-visual models have also experienced rapid development in recent yearsradford2021learning ; lu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . The diverse prior assumption associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others. If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, it may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration. On the contrary, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing the cohesion and autonomy becomes progressively more challenging in a nonlinear manner.
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
Figure 1: Language as Reference Framework (LaRF)
Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, AllSpark, integrating thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data.
To maintain the autonomy between modalities, AllSpark introduce specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, there exists significant dimensional gap between the tokens of each modality and the language modality. To address this issue, we introduce modality bridge, a mechanism from Lynxzeng2023matters to accomplish the dimensional mapping from each modality’s tokens to the language tokens.
Finally, considering the existing gap between the interpretation by the multimodal large language model and the specific downstream tasks, we have designed task heads for each downstream task to enhance the model’s generalization capability. It is noteworthy that, given the powerful interpretation capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads
Experiments demonstrate that AllSpark, without the expert knowledge in most modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatiotemporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, AllSpark’s accuracy is only 0.84 lower than the optimal model, and in the trajectory modality, the ADE metric differs by only 0.07 compared to the optimal model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
In a word, our contributions can be summarized as follows:
  • 1.
    To our best knowledge, we first propose a unified multimodal spatiotemporal general model, AllSpark, successfully integrating 13 spatiotemporal modalities into a single model.
  • 2.
    Inspired by the human cognitive system and linguistic philosophy, we propose the Language as Reference Framework (LaRF), offering a novel solution to balance cohesion and autonomy among multiple modalities. This principle provides a fresh approach to the multimodal unification.
  • 3.
    We provide new possibilities for the paradigm of designing multimodal models in the spatiotemporal domain. MSTAGI demonstrates the potential and capabilities of building general model with large language models, thereby contributing to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
  • 4.
    Experiments indicate that AllSpark, without expert knowledge of the majority of spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
2 Related Work
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatiotemporal domain is the continual increase in the number of modalities for joint interpretation. Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to build multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult. In the following, we recall the development of intelligent methods in the spatiotemporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
2.1 Single-Modal Model
For one-dimensional modalities, we focus on Code and Text. Code and Text are fundamentally language modalities, but due to the excellent characteristics of Code, such as strict syntax, unambiguous nature, and its ability to interact with machines, we treat it as a separate modality. Z. Feng et al. have pre-trained a model, CodeBERTfeng2020codebert , which facilitates the mutual transfer of information between Code and natural language modalities. For the Text modality, landmark contributions include Transformervaswani2017attention ,BERTkenton2019bert ,GPT series radford2018improving ; radford2019language ; brown2020language ; openai2023gpt4 , which have inspired a subsequent series of works.
We categorize RGB, multispectral, hyperspectral, infrared, SAR, Table, Graph, Trajectory, and Oblique Photography as two-dimensional modalities.
Among the two-dimensional modalities, the standard three-channel RGB image is one of the most common. For the RGB modality, the ResNethe2016deep proposed by K. He et al., based on the importance of visual global and local information, and the ViTdosovitskiy2020image introduced by A. Dosovitskiy et al., leveraging a global attention mechanism, represent two landmark contributions.
An increase in the number of channels in images leads to the multispectral and hyperspectral modalities. B. Huang et al. proposed the STDCNNhuang2018urban , leveraging the characteristic of a higher number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images. In comparison to multispectral images, hyperspectral images have even more bands, often reaching hundreds. Based on this, X. Yang et al. introduced the R-3D-CNNyang2018hyperspectral to further enhance the extraction of spectral features.
Beyond an increase in the number of bands, attention has been directed towards images captured in specific spectral bands. For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention. H. Chen et al. proposed a shallow convolutional neural network, integrating a decision tree, to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoringchen2020deep . In the case of SAR images, formed by active microwave radar, S. Chen et al. introduced AConvNetchen2016target , a widely used fully convolutional neural network for SAR image intelligent interpretation.
For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from different angles to reconstruct a three-dimensional model. Y. Yao et al. addressed this task by proposing MVSNetyao2018mvsnet . Utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing the three-dimensional model of the target from the features extracted from two-dimensional images.
Beyond the visual domain, Table is one of the commonly used modalities for recording and expressing information. TabNetarik2021tabnet , proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
The trajectory modality reflects the temporal changes in the spatial positions of objects. A. Gupta et al. proposed Social GANgupta2018social based on the characteristic of trajectory multiplicity. This model combines historical trajectory information with social context to predict multiple plausible future outcomes.
For the Graph modality, Kipf and Welling introduced the classic Graph Convolutional Network (GCN)kipf2016semi , which is based on the adjacency relationships between nodes in the graph. P. Velicˇkovic et al. proposed GATvelivckovic2017graph , which incorporates attention mechanisms into the Graph modality.
Finally, we turn our attention to the three-dimensional modalities: Video and Point Cloud.
The Video modality is formed by stacking images over time. A model proposed by A. Karpathy et al. extends the connectivity of CNN models over time. Additionally, it utilizes multi-resolution techniques to accelerate the training process, effectively reducing the impact of spatiotemporal redundancy on model efficiency in the Video modalitykarpathy2014large .
The Point Cloud modality captures information about the position, shape, color, texture, and other aspects of three-dimensional objects. C. R. Qi et al. introduced the classic PointNetqi2017pointnet for the Point Cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features. W. Wu et al. extended convolution operations to three-dimensional point clouds with the introduction of PointConvwu2019pointconv .
2.2 Multi-modal Model
While the single-modal methods in Section 2.1 have demonstrated excellent performance within their respective modalities, they often face challenges in generalizing across multiple modalities due to their construction based on specific prior assumptions. Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavored to balance the cohesion and autonomy among modalities to build multimodal models.
For joint interpretation of two modalities, H. Li and X.-J. Wu recognized the enhancement possible through the fusion of RGB and Infrared images, which convey information about color, texture, and temperature. Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and Infrared images: DenseFuseli2018densefuse . This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of RGB and Infrared modalities. A. Sadeghian et al. extended the use of historical trajectories for predicting multiple plausible trajectories in Social GANgupta2018social . They introduced RGB images to enhance scene data in Sophiesadeghian2019sophie , achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR imageshughes2020deep . X. Li et al. introduced DTCDNli2021deep , a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral imagesyang2018hyperspectralandmultispectral .
Notably, CLIPradford2021learning proposed by A. Radford et al. associates the RGB modality with the text modality using contrastive learning. With pretraining guided by weak supervisory signals from text, CLIP demonstrates outstanding capabilities in both visual single-modal tasks and visual-text multimodal tasks, inspiring a series of subsequent workslu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . Y. Zhang et al. introduced Meta-transformerzhang2023meta , leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities such as point clouds, infrared, and hyperspectral. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, successfully unifying eight modalities including Image, Audio, Video, and Pointhan2023onellm . The success of these approaches imply the unique role of language modalities in multimodal model.
Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle, Language as Reference Framework (LaRF). Guided by this principle, we balance the cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies 13 spatiotemporal modalities and possesses the potential to extend to an arbitrary number of modalities.
3 Method
3.1 Theory
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
Inspired by our observation, to balance the cohesion and autonomy among multiple modalities, we introduce the fundamental principle, Language as Reference Framework (LaRF). The significance is:
  1. 1.
    Alignment Capability: Language can accurately encode both cohesion and autonomy information across multiple modalities. Aligning each modality with the language modality enables the attainment of a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
  2. 2.
    Reasoning Capability: Language, being a tool for human thought and expression, inherently possesses the ability to perform complex reasoning. Each modality, when represented in a unified space with LaRF, inherits the reasoning capability of language, unlocking the potential for multimodal joint reasoning.
  3. 3.
    Interpretability: Deep learning have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
  4. 4.
    Interactivity: Language not only aids humans in understanding intelligent models but also facilitates intelligent models in understanding humans. In an intelligent system guided by the LaRF, humans can directly express their needs using natural language. This iterative correction of the model’s output based on human interaction will become a new paradigm for the training and inference of intelligent models.
  5. 5.
    Scalability: The multimodal system guided by the LaRF is agnostic to specific modalities. New modalities only need to establish a mapping to the language model to participate in joint reasoning with other modalities. Therefore, theoretically, a multimodal model based on LaRF can be extended to an arbitrary number of modalities.
In summary, we contend that intelligent models, when engaged in multimodal perception processes akin to those of humans, should adopt language as the reference framework for abstract concepts across modalities. The analysis and reasoning should occur in the feature space of language. As illustrated in the conceptual model in Figure 2, when constructing specific multimodal models, data from each modality should be encoded into modality-independent tokens with contextual information. These modality-specific tokens should align with the tokens of language. A universal intelligent model, primarily based on the Transformer architecture, can then be employed for parsing and reasoning, ultimately generating a comprehensive multimodal representation.
Figure 2: The conceptual model with LaRF
3.2 Overview
Table 1: List of symbols
Module Variants
Symbol
Meaning
Symbol
Meaning
$M$
AllSpark
$m_{i}$
modal i
$f_{i}$ modal encoder for $q$ query vectors of the bridge
$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
T text tokenizer $p_{i}$ text prompt for $m_{i}$
$F$
Multi-modal LLM
$\theta$
model parameters
$H_{task}$
Task Head
$y$
label
$L$ loss function $t_{i}$ tokens of $m_{i}$
${Enc}_{n}$
n-layer transformer encoders
$W$
weights of the linear layer
$Emb$
embedding layer
$w_{i}$
word of the code or text modal
${Enc}_{resnet}$
resnet
$Sparse$
sparse module of the TubeViT
$Grouper$
point grouper of the PointBERT
$Conv1d$
1D convolution layer
$FFN$
feed forward network
Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: Modal Encoder, Modal Bridge, Text Tokenizer, Multimodal Large Language Model, and Task Head. The overall architecture is depicted in Figure 3.
Figure 3: AllSpark Architecture
To maintain the autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (Section 3.3 for details). However, the dimens ion of tokens outputted by modal encoders are still inconsistent. To achieve the goal of parsing in the unified representational space of language, we introduced the modal bridge from the Lynxzeng2023matters . The modal bridge aims to project tokens from each modality into the dimension of the Multimodal Large Language Model (Section 3.4 for details). The formalization of this process is as follows:
$s_{i}=\Phi\left(f_{i}\left(m_{i}\right),q\right)$ (1)
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the Multimodal Large Language Model. Each modality’s input data $m_{i}$ is mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table 1 summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified Multimodal Large Language Model to parse data from various modalities. The text tokenizer and Multimodal Large Language Model in AllSpark are based on the visual-language model Lynxzeng2023matters . To extend Lynx to 13 spatiotemporal modalities, we design specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatiotemporal modalities. Finally, we acknowledge that there exists a gap between the parsing results of the model and the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
The entire model M can be formalized as:
$M\left(m_{i},p_{i}\right)=H_{task}\left(F\left(s_{i}\oplus T\left(p_{i}\right)\right)\right)$ (2)
Where $p_{i}$ represents the text prompt of $m_{i}$, $T$ denotes the text tokenizer, $\oplus$ indicates the concatenation operation of text tokens and modality tokens in the sequence, $F$ represents the Multimodal Large Language Model, and $H_{task}$ signifies the task head.
All tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model. The optimization objective of the model can be formalized as:
$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right),\theta_{i}\right)}$ (3)
3.3 Independent Encoder for Each Modality
The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as: $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$. We have designed different modal encoders for each modality to maintain autonomy among modalities. The following provides individual introductions for each modality:
3.3.1 1D modal
Code & Text: Code is essentially a specialized form of language, but due to its distinct properties—such as having a strict syntax, being unambiguous, and being capable of interacting with machines—we separate it as a distinct modality. Exploring intelligent methods for the Code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the Code and Text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in text or code.
3.3.2 2D modal
RGB: RGB imagery represents the visible light spectrum, reflecting the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality in visual research. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA. EVA is a large visual backbone model composed of 40 stacked Transformer Blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them for training.
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. Typically, it includes more than the three visible light bands present in the RGB modality. MSI incorporates multiple non-visible light bands such as near-infrared, short-wave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard Vision Transformerdosovitskiy2020image , modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information about the terrain. Unlike RGB and multispectral imagery, where an image serves as a single sample, in hyperspectral imagery, all bands for each pixel are treated as a single sample, i.e., $m_{HSI}\in R^{1*1*C}$. In the modal encoder for hyperspectral imagery, we first use a linear projection layer to expand the feature dimensions for each pixel. This process is formalized as $W*m_{HSI}^{T}$, where $W\in R^{1*d}$ is the weight matrix of the linear projection layer. Subsequently, we use a 12-layer standard Transformer encoder to extract its features. The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
Table: The Table modality consists of several rows and columns, where each row represents a record, and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the Table modality inherits the design from TabFormerpadhi2021tabular : first, based on the different degrees of discreteness for each column attribute, we use independent Embedding Layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer Transformer Encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
Trajectory: The trajectory modality reflects the changing information of an object over time and space, composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points. The encoder for the trajectory modality inherits the design from TUTRshi2023trajectory : first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features. This step can be formalized as ${W*m}_{Trajectory}$, where $W\in R^{d*\left(l*2\right)}$ is the weight matrix of the linear projection layer. Then, we use a 2-layer Transformer Encoder to extract its features. The entire process can be formalized as $f\left(m_{Trajectory}\right)={Enc}_{2}\left({W*m}_{Trajectory}\right)$.
SAR: The Synthetic Aperture Radar (SAR) modality is a type of active remote sensing that reflects the electromagnetic characteristics of objects with respect to microwave backscatter. Due to different polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$. Therefore, we designed a simple three-layer convolutional network as its modal encoder.
Infrared: The dataset we selected for infrared is SYSU-MM01, which is, in fact, a multi-modal re-identification dataset comprising both visible and near-infrared modalities. The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$. We inherited and slightly modified the design of DDAGye2020dynamic , discarding the shared ResNet-50 layers’ parameters in DDAG and employing two independently parameterized ResNet-50 networks. These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{resnet}\left(I_{r}\right)\oplus\ {Enc}_{resnet}\left(I_{i}\right)$.
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the graph’s features, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes, and d is the node feature dimension. In AllSpark, the graph’s modality encoder is based on STAEformerliu2023spatio , with the main design idea being to first use a linear layer to extend its feature dimension and then use several Embedding Layers to separately encode features such as the node’s own characteristics, spatial characteristics, and temporal characteristics in the graph. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{video}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view, and n is the number of views. In the encoder for oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them. The process can be formalized as $f\left(m_{video}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
3.3.3 3D modal
Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$. The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions. To address this issue, the encoder for the video modality in AllSpark is based on TubeViTpiergiovanni2023rethinking , which performs sparse sampling on the highly redundant data in both time and space to reduce redundancy. The features obtained from sparse sampling are then fed into a 6-layer Transformer Encoder for further feature extraction. The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left(Sparse(m_{video})\right)$.
PointCloud: PointCloud are typically composed of three-dimensional coordinates and their feature values, $m_{pointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points, and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, color, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERTyu2022point : firstly, point cloud data is grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{pointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups, and $N$ represents the number of points in each group. Next, the grouped results are inputted into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are inputted into a standard 12-layer Transformer Encoder to extract their global features. The entire process can be formalized as $f\left(m_{pointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{pointCloud}))\right)$.
3.4 Modal2Language Bridge
The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model, and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities. The keys and values in the cross-attention layer are the features outputted by the modal encoders.
The feedforward neural network inherits the classic design from the original Transformer, consisting of two linear layers with an inserted activation layer. The entire process can be formalized as:
$\Phi\left(Q,s_{i}\right)={FFN}\left({softmax}\left(\frac{W_{q}Q^{T}\left(w_{k}s_{i}^{T}\right)^{T}}{\sqrt{D}}\right)*(W_{v}s_{i}^{T})\right)$ (4)
Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$ respectively.
3.5 Task-Guided Text Prompts and Task-Heads
To extend the visual-language multimodal model to 13 spatiotemporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table 2 provides a list of all the text prompts.
Table 2: List of prompts
Dimension
Modal
Prompt
1D
Text
1.Please determine if this movie review is positive or negative?
Code
2D
RGB
1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
MSI
1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
HSI
1.Given the spectral information, can you help determine which class this pixel belongs to? 2.Here is the spectral data for a pixel. Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel? 3.The spectral information for a pixel is given, but the data is noisy. Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
Table
1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
Oblique Photography
1.Please reconstruct the 3D model from the given images of five different views.
Trajectory
1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian. 2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian. 3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
SAR
1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
Infrared
1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape. Then, locate the person with matching features in a series of infrared images
Graph
1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week. Identify any patterns or trends and predict the traffic conditions
3D
Video
1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
Point Cloud
1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to. Focus on the shape and structure evident in the point cloud
Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks. To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task. Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible. The detailed summary of each modality’s task heads can be found in Table 4.
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, Video, and PointCloud modalities. Although the task on the HSI modality is a segmentation task, it can be formalized as a per-pixel classification task. Similarly, for the re-identification task on the Infrared modality, it can also be formalized as a classification task. Therefore, a single-layer linear layer is used as task head for the mentioned modalities.
For regression tasks on the Table, Trajectory, and Graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx model itself is a language model, the Code and Text modalities directly use its native text decoder. The task head for the Oblique Photography modality is based on the 3D reconstruction architecture from the Ada-MVSliu2023deep model.
4 Experiment
4.1 Setup
Our experiments aimed to evaluate the adaptability of AllSpark to various spatiotemporal modalities. To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain. The specific details on dataset selection refer to Section 4.2.
Additionally, for result reproducibility, we employed similar experimental settings across all modalities. We used the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments. Table 3 summarizes the hyperparameter settings for experiments on each modality.
Table 3: Summary of hyperparameters
Dimension
Modal
Max Lr
Max Epochs
Warm-up Epochs
1D
Text
9.0*10-6
5 1
Code
1.0*10-5
4 1
2D
RGB
5.0*10-5
50 5
MSI
2.0*10-5
50 5
HSI
1.0*10-4
30 3
Infrared
5.0*10-5
50 5
SAR
9.0*10-6
30 3
ObliquePhotography 5.0*10-5 30 3
Table
2.0*10-5
30 3
Trajectory
1.0*10-5
30 5
Graph
8.0*10-5
10 2
3D
PointCloud
3.0*10-5
100 10
Video
1.0*10-5
3 1
4.2 Dataset
Table 4 summarizes the data and tasks used in experiments for each modality. Below, we will provide detailed explanations in order:
Text: The IMDB dataset is a binary sentiment analysis dataset, consisting of 50,000 reviews from the Internet Movie Database (IMDb) labeled as positive or negative. Additionally, the dataset includes some unlabeled data. In our experiments, only the labeled data from the IMDB dataset was utilized for supervised sentiment classification tasks.
Code: CodeSearchNet is a large-scale dataset of function code and its documentation from GitHub, covering six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby. The task performed on the code modality is Code Document Generation, and we tested it on the Ruby and JavaScript languages.
RGB: NWPU-RESISC45 is a large-scale open dataset for visible light remote sensing image scene classification. It includes 45 land-use categories such as airplane, baseball diamond, beach, commercial area, etc. Each category has 700 remote sensing images, making a total of 31,500 images in the dataset. The image size is 256*256 pixels, and we selected a version of the dataset split by the official release, using 20% of the data for training.
MSI: The EuroSAT dataset is a multispectral modality dataset for land use and land cover (LULC) classification. The samples are sourced from the Sentinel-2 optical satellite and include all 13 bands. The data is categorized into 10 classes, with a total of 27,000 images. We adopted a random 9:1 split for training and testing.
HSI: The Pavia University dataset is a high-spatial-resolution hyperspectral dataset acquired by the ROSIS sensor. It comprises 103 bands with a size of 610*340 pixels. The dataset includes nine land cover categories such as Asphalt, Meadows, Gravel, etc. We used a 4:6 split for training and testing.
Table: The PRSA dataset is a collection of air quality data from multiple stations in Beijing, containing hourly measurements of air pollutants. The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations. In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, wind direction). The task was to predict the concentration of PM2.5. We split the data into training (40%) and testing sets.
Trajectory: The ETH-UCY dataset is one of the widely used benchmarks for pedestrian trajectory prediction, divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2. In our experiments, we utilized the ETH subset.
SAR: The MSTAR dataset is a synthetic aperture radar dataset designed for military stationary target recognition, comprising ten categories of military targets. We employed the standard operating conditions (SOC) dataset preprocessing method proposed by S. Chen et al. chen2016target , ensuring that the serial numbers and target configurations are consistent between the test and training sets, while the aspects and depression angles differ.
Infrared: SYSU-MM01 is a cross-modal human re-identification dataset with optical-infrared modalities, including images from 4 RGB cameras and 2 infrared cameras of 491 different individuals. We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals, while the test set consists of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for re-identification targets.
Graph: METR-LA is a traffic data collection from loop detectors on the highways of Los Angeles, spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
Oblique Photography: WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, we use area 1 as the training set and area 2 as the test set.
Video: UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels, sourced from YouTube.
PointCloud: ModelNet40 is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects. We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
4.3 Results
4.3.1 RGB
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the Top-1 Accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore in Table 4, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
Table 4: RGB scene classification with AllSpark(80% as trainset)
Method
Publication
Acc (%)
CNN-CapsNetzhang2019remote
RS2019
89.03
DFAGCNxu2021deep
TNNLS2021
89.29
D-CNN with GoogleNetcheng2018deep
TGRS2018
90.49
D-CNN with VGGNetcheng2018deep
TGRS2018
91.89
SCCovhe2019skip
TNNLS2019
92.1
SeCo-ResNet-50manas2021seasonal
ICCV2021
92.91
MG-CAPwang2020multi
TIP2020
92.95
LSENetbi2021local
TIP2021
93.34
MSANetzhang2021multiscale
JSTARS2021
93.52
IDCCPwang2020invariant
TGRS2021
93.76
MBLANetchen2021remote
TIP2021
94.66
GRMANet-ResNet-50li2021gated
TGRS2021
94.72
EMSNetzhao2023emscnet
TGRS2023
95.37
ViTAE-B + RVSAwang2022advancing
TGRS2022
95.69
AllSpark
ours
94.85
4.3.2 MSI
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was Top-1 Accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups in Table 5: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
Table 5: MSI Land Cover Classification with AllSpark
Expert Knowledge
Method
Publication
Accuracy (%)
No
ResNet-18manas2021seasonal
ICCV2021
63.21
ResNet-50nampally2023performance
IGARSS2023
91.13
InceptionNetzhang2020semi
ICIP2020
93.07
EfficientNetgomez2021msmatch
IJSTARS2021
93.94
AllSpark ours 94.03
ResNet-152nampally2023performance
IGARSS2023
96.63
Yes
MoCoV2manas2021seasonal
ICCV2021
89.51
SeComanas2021seasonal
ICCV2021
93.14
SEERgoyal2022vision
Arxiv2022
97.6
DINO-MCwanyan2023dino
Arxiv2023
98.78
4.3.3 HSI
We conducted pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include Overall Accuracy (OA), Average Accuracy (AA), and Kappa. Since AllSpark does not possess expert knowledge in the hyperspectral modality, we compared it with the semi-supervised baselines summarized by D. Uchaev and D. Uchaev uchaev2023small . The results in Table 6 demonstrate that AllSpark outperforms many hyperspectral image classification methods such as IFRF and S-DMM, with a difference of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
Table 6: HSI pixel Classification with AllSpark
Method
OA
AA
Kappa
3D-CNN
75.24 80.26 68.34
CA-GAN
76.81 76.94 71.02
3D VS-CNN
81.63 83.86 76.46
RPNet
84.92 83.26 80.52
S-DMM
88.3 93.76 84.9
IFRF
88.38 85.99 84.97
AllSpark 89.18 86.65 85.32
DCFSL
90.71 90.2 87.73
TC-GAN
93.2 91.6 91
PRNet-RF
95.6 94.96 94.27
4.3.4 Table
Table 7: PM2.5 prediction with AllSpark
Expert Architecture Method PM2.5
RMSE MAE R2
No
GWOxie2020evolving
62.2 40.8 -
TabBERTpadhi2021tabular
32.8 - -
AllSpark 29.03 18.04 0.87
Yes
Stacked ResNet-LSTMcheng2022stacked
40.68 23.75 0.8
CBAM-CNN-Bi LSTMli2022prediction
18.9 11.2 0.94
For the Table modality, we evaluated AllSpark on the regression prediction task using the PRSAzhang2017cautionary dataset. The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section 4.2). The performance metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared (R2).
Table 7 presents the comparative results between AllSpark and baselines. It’s worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi LSTM proposed by D. Li et al.li2022prediction uses CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the Stacked ResNet-LSTM model proposed by X. Cheng et al.cheng2022stacked employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: with expert structures and without expert structures. In models without expert structures, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method in models with expert structures (CBAM-CNN-Bi LSTM). This reflects the excellent adaptability of AllSpark to the Table modality.
4.3.5 Code
Table 8: Code Document Generation with AllSpark
Expert Knowledge Method MRR
Ruby Javascript
No
BIRNN[68]
0.084 0.153
1D-CNN[68]
0.245 0.352
selfAtt[68]
0.365 0.451
NBoW[68]
0.429 0.461
RoBERTa[67]
0.625 0.606
AllSpark 0.627 0.635
Yes
RoBERTa(Code)[67]
0.661 0.64
CodeBERT[67]
0.693 0.706
GraphCodeBERT[67]
0.732 0.711
For code modality, we evaluated the performance of AllSpark on the Code Document Generation task using the CodeSearchNet dataset. This task involves generating corresponding documents based on the provided function code. We conducted tests for both Ruby and JavaScript languages, using Mean Reciprocal Rank (MRR)guo2020graphcodebert as the evaluation metric.
Similar to the previous modalities, since MSI-AGI does not possess expert knowledge in the Code modality, we categorized the baselines into two groups: with and without expert knowledge. As shown in Table 8, AllSpark achieved the SOTA results in the group without expert knowledge and remained comparable to models with expert knowledge. This demonstrates the strong adaptability of AllSpark to the Code modality.
4.3.6 PointCloud
For pointcloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with Top-1 Accuracy as the metric. In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties like permutation invariance and symmetry in 3D point clouds. However, these structures designed for the unique priors of the modality are challenging to transfer across modalities. Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
Table 9: PointCloud classification with AllSpark
Method
Expert Architecture Expert Knowledge Acc(%)
PointNetqi2017pointnet
Yes No 89.2
Kd-netklokov2017escape
Yes No 90.6
SPH3D-GCNlei2020spherical
Yes No 91.4
PointNet++qi2017pointnet++
Yes No 91.9
SO-Netli2018so
Yes No 92.5
PointVGGli2021pointvgg
Yes No 93.6
PointBERTyu2022point
No Yes 93.8
PointGPTchen2023pointgpt
Yes Yes 94.9
AllSpark
No No 91.2
The design of AllSpark is based on a general sequence-to-sequence architecture. As shown in Table 9, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-net). It maintains comparability with the state-of-the-art model, PointGPT, which has both a modal expert structure and modal expert knowledge. This suggests that AllSpark holds significant potential for applications in the point cloud modality.
4.3.7 Trajectory
Table 10: Trajectory predicition with AllSpark
Method
ADE/FDE
Social GAN
0.87/1.62
SoPhie
0.70/1.43
STAR
0.36/0.64
SGCN
0.63/1.03
CAGN
0.41/0.65
SIT
0.39/0.62
SocialVAE
0.47/0.76
PCENet
0.54/0.87
AgentFormer
0.45/0.75
MemoNet
0.40/0.61
SocialVAE+FPC
0.41/0.58
TUTR
0.40/0.61
AllSpark
0.43/0.69
For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task. This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period. We report accuracy using Average Displacement Error (ADE) and Final Displacement Error (FDE)shi2023trajectory . Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (Ground Truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, ADE and FDE are used to measure their $L2$ distances, calculated as follows:
${ADE}=\frac{1}{T_{pred}}\sum_{t=T_{bos}}^{T}\sqrt{\left(x_{t}-\widehat{x_{t}}\right)^{2}+\left(y_{t}-\widehat{y_{t}}\right)^{2}}$ (5)
${FDE}=\sqrt{\left(x_{T}-\widehat{x_{T}}\right)^{2}+\left(y_{T}-\widehat{y_{T}}\right)^{2}}$ (6)
In Table 10, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, using a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to the SOTA model (STAR) with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This suggests that AllSpark demonstrates excellent adaptability to the trajectory modality.
4.3.8 SAR
Table 11: SAR classification with AllSpark
Method
Accuracy (%)
EMACHsrinivas2014sar
88
SVMsrinivas2014sar
90
AdaBoostsrinivas2014sar
92
MSRCdong2014sparse
93.6
IGTsrinivas2014sar
95
MSSdong2015classification
96.6
Cond Gausso2001sar
97
M-PMCpark2014modified
98.8
AConvNetschen2016target
99.13
AllSpark
88
For SAR modality, AllSpark’s adaptability was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is Top1 Accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC (standard operating conditions) settings from AConvNetschen2016target . Table 11 presents the comparison results between AllSpark and the state-of-the-art model under these settings. Since AllSpark lacks SAR image modality expert knowledge and the spatial invariance prior to changes in perspective, its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%. This is comparable to the EMACH method. Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for our future research.
4.3.9 Infrared
AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human re-identification task involving RGB and infrared modalities. In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images. Table 12 presents the comparison results between AllSpark and the state-of-the-art re-identification methods, with the evaluation metric being Rank at 20 accuracy. The results indicate that AllSpark (76.31%) achieved similar accuracy compared to HSME (77.95%). While our method did not reach the highest accuracy, it has already surpassed some expert models in RGB-Infrared re-identification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, showcasing the potential of AllSpark in handling the infrared modality.
Table 12: Visible-Infrared Re-identification with AllSpark
Method
Rank-20(All-search)
Two-streamwu2017rgb
65.5
One-streamwu2017rgb
66.74
TONEye2018hierarchical
68.6
HCMLye2018hierarchical
69.17
Zero-Padwu2017rgb
71.33
AllSpark 76.31
HSMEhao2019hsme
77.95
BDTRye2019bi
81.07
DDAGye2020dynamic
95.81
4.3.10 Graph
For the Graph modality, AllSpark’s performance was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include RMSE, MAE, and R2. Table 13 presents the comparison between AllSpark and the state-of-the-art results on the METR-LA dataset, with the baseline derived fromliu2023spatio . The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of RMSE, demonstrating AllSpark’s excellent adaptability to the Graph modality.
Table 13: Traffic prediction with AllSpark
Method
RMSE MAE R2
HI
6.8 14.2 10.15
GWNet
3.51 7.28 9.96
DCRNN
3.54 7.47 10.32
AGCRN
3.59 7.45 10.47
STGCN
3.6 7.43 10.35
GTS
3.59 7.44 10.25
MTGNN
3.47 7.21 9.7
STNorm
3.57 7.51 10.24
GMAN
3.44 7.35 10.07
PDFormer
3.62 7.47 10.91
STID
3.55 7.55 10.95
STAEformer
3.34 7.02 9.7
AllSpark
3.81 7.52 11.24
4.3.11 Language
Table 14: Text understanding with AllSpark
Method
Accuracy (%)
RoBERTaliu2019roberta
95.3
ULMFiThoward2018universal
95.4
BERTxie2020unsupervised
95.49
Mixed VATsachan2019revisiting ; miyato2016adversarial
95.68
LongFormerbeltagy2020longformer
95.7
XLNetyang2019xlnet
96.8
ERNIE-Doc-Largeding2020ernie
97.1
AllSpark
96.78
We tested AllSpark’s natural language processing capabilities on the IMDB datasetmaas2011learning with the task of binary sentiment classification (positive or negative). AllSpark loaded weights from the Lynx, and therefore, it can be considered to possess expert knowledge in the natural language modality. When compared with the state-of-the-art models on the IMDB dataset, as shown in Table 14, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
4.3.12 Video
For the Video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset. The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being Top1 accuracy. In Table 15, we compared AllSpark with the current state-of-the-art models. Currently, AllSpark’s adaptation to the Video modality is not optimal, showing a significant gap compared to baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark, and we trained it for only 3 epochs on the UCF101 dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective in capturing temporal information.
Table 15: Video classification with AllSpark
Method
Accuracy (%)
OPN
59.6
VCOP
72.4
SpeedNet
81.1
VTHCL
82.1
CVRL
94.4
VideoMAE v1
96.1
VideoMAE v2
99.6
AllSpark
27.5
4.3.13 Oblique Photography
For the Oblique Photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset. The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used is the Percentage of Accurate Grids in Total (PAG)liu2023deep , calculated by the following formula:
$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
The suffixes in PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 meters, and ${PAG}_{10}$ indicates an error within 1 meter. Table 16 presents the comparison of AllSpark with popular multi-view 3D reconstruction models on the WHU-OMVS dataset. AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, it lacks the ability to design specific model structures for processing, severely restricting its performance. Adapting multi-modal large models to dense prediction tasks and optimizing the architecture is one of our future research directions.
Table 16: 3D reconstruction with AllSpark
Method
${PAG}_{6}$
${PAG}_{10}$
MVSNet
81.15 91.44
CasMVSNet
95.45 98.02
Ada-MVS
96.14 98.1
UCSNet
96.25 98.45
AllSpark
6.4 10.4
5 Discussion
5.1 Limitation
Certainly, our work has some limitations, and these will guide our future research directions:
  1. 1.
    Lack of interaction between different modalities. While AllSpark incorporates interaction between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks. This is primarily due to constraints such as the lack of multi-modal paired data, the inflexibility of the model architecture, and the limited richness of multi-modal interaction tasks. In the future, we plan to explore and work more extensively in these directions.
  2. 2.
    Lack of in-depth research on multi-modal alignment methods. While we introduced the basic principles of the Language as Reference Framework (LaRF) and aligned all modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method. We haven’t conducted in-depth research into the various methods of multi-modal alignment. Moreover, our testing focused on the model’s adaptability to different modalities, and we haven’t explored potential capabilities such as interactive reasoning, zero-shot learning, and multi-modal emergence that the model may possess after alignment with language. In the future, we plan to delve deeper into researching multi-modal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
  3. 3.
    Initial work. Our current work is still in its initial exploratory phase, and we haven’t carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, SAR, and video. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
  4. 4.
    Expensive Cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. Although we freeze most parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities remains one of our future research directions.
  5. 5.
    Interesting phenomenon. During our exploratory experiments, we discovered some interesting properties of large models. For instance, spatial information tends to degrade in large models, leading to collapse when performing dense prediction tasks such as segmentation and detection. Additionally, these models struggle to optimize when transferred to a small amount of downstream data. Large models often require smaller hyperparameters and are sensitive to them. These observations might partially reveal the working mechanisms of large models, and we plan to conduct more in-depth investigations into these phenomena in the future.
5.2 Potential of the LaRF
Inspired by the human cognitive system and linguistic philosophy, we propose ”Language as Reference Framework (LaRF)” as the first principle for constructing our unified multimodal model. Its foreseeable potential includes at least the following three points:
  1. 1.
    Efficient Generalization of Large Models. Currently, the computational power and data scale required for training large models are rapidly expanding, and even the cost of fine-tuning large models is becoming prohibitive. Therefore, in the future, training large models for every domain is almost an impossible task. Language, however, holds the potential for achieving efficient generalization of large models, expanding them from their native domains to additional domains at minimal cost. In our proposed AllSpark, we have designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion. In theory, AllSpark can extend to any number of modalities. In the future, we will conduct more in-depth research on the impact of text prompts and lightweight parameter modules on the generalization of large models.
  2. 2.
    Interpretable Reasoning. Deep learning models have often been referred to as ”black boxes,” indicating that the reasoning process of these models is invisible and challenging to interpret. Some research on the interpretability of deep learning models is often based on complex mathematical models and numerous assumptions, greatly limiting the practical applications of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality is crucial. However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to directly use natural language for outputting reasoning chains and justifications.
  3. 3.
    Transition from end-to-end to interactive paradigm. The end-to-end paradigm refers to the learning approach where the model takes input and directly outputs results. In recent years, the end-to-end paradigm has become increasingly popular due to its simple and clear architecture and excellent performance. However, it also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues. A LaRF-based architecture has the potential to achieve an interactive paradigm, where users input raw data and corresponding text prompts, and the model automatically performs relevant operations based on the prompts. Users can even iteratively adjust the text prompts based on the results. Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to the end-to-end paradigm.
6 Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models should follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing 13 spatiotemporal modalities. Experimental results demonstrate that AllSpark exhibits excellent adaptability and application potential across various spatiotemporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into mechanisms guided by natural language, efficient generalization of large models, and the transition to an interactive paradigm.